# The ELM Neuron: an Efficient and Expressive

### v1

@@ Abstract
Traditional large-scale neuroscience models and machine learning utilize simplified models of individual neurons, relying on collective activity and properly adjusted connections to perform complex computations. However, each biological cortical neuron is inherently a sophisticated computational device, as corroborated in a recent study where it took a deep artificial neural network with millions of parameters to replicate the input-output relationship of a detailed biophysical model of a cortical pyramidal neuron. We question the necessity for these many parameters and introduce the Expressive Leaky Memory (ELM) neuron, a biologically inspired, computationally expressive, yet efficient model of a cortical neuron. Remarkably, our ELM neuron requires only 8K trainable parameters to match the aforementioned input-output relationship accurately. We find that an accurate model necessitates multiple memory-like hidden states and intricate nonlinear synaptic integration. To assess the computational ramifications of this design, we evaluate the ELM neuron on various tasks with demanding temporal structures, including a sequential version of the CIFAR-10 classification task, the challenging Pathfinder-X task, and a new dataset based on the Spiking Heidelberg Digits dataset. Our ELM neuron outperforms most transformer-based models on the Pathfinder-X task with $77\%$ accuracy, demonstrates competitive performance on Sequential CIFAR-10, and superior performance compared to classic LSTM models on the variant of the Spiking Heidelberg Digits dataset. These findings indicate a potential for biologically motivated, computationally efficient neuronal models to enhance performance in challenging machine learning tasks.
@@ 1 Introduction
The human brain has impressive computational capabilities, yet the precise mechanisms underpinning them remain largely undetermined. Two complementary directions are pursued in search of mechanisms for brain computations. On the one hand, many researchers investigate how these capabilities could arise from the collective activity of neurons connected into a complex network structure Maass (1997); Gerstner and Kistler (2002); Grüning and Bohte (2014), where individual neurons might be as basic as leaky integrators or ReLU neurons. On the other hand, it has been proposed that the intrinsic computational power possessed by individual neurons Koch (1997); Koch and Segev (2000); Silver (2010) contributes a significant part to the computations.
@@
Even though most work focuses on the former hypothesis, an increasing amount of evidence indicates that cortical neurons are remarkably sophisticated Silver (2010); Almog and Korngreen (2016); Gidon et al. (2020); Larkum (2022), even comparable to expressive multilayered artificial neural networks Poirazi et al. (2003); Jadi et al. (2014); Ujfalussy et al. (2018); Tzilivaki et al. (2019); Beniaguev et al. (2021); Marino (2021); Jones and Kording (2021, 2022); Poirazi and Papoutsi (2020), and capable of discriminating between dozens to hundreds of input patterns Poirazi et al. (2003); Gütig and Sompolinsky (2006); Hawkins and Ahmad (2016); Moldwin and Segev (2020). Numerous biological mechanisms, such as complex ion channel dynamics (e.g. NMDA nonlinearity Major et al. (2013); Lafourcade et al. (2022); Tang et al. (2023)), plasticity on various and especially longer timescales (e.g. slow spike frequency adaptation Kobayashi et al. (2009); Bellec et al. (2020)), the intricate cell morphology (e.g. nonlinear integration by dendritic tree Stuart and Spruston (2015); Poirazi and Papoutsi (2020); Larkum (2022)), and their interactions, have been identified to contribute to their complexity.
Thus, one of the central questions in computational neuroscience is which learning and memory capabilities can be attributed to an individual cortical neuron and how mechanisms within a single neuron can work together to generate the observed computational complexity.
@@
Detailed biophysical models of cortical neurons aim to capture this inherent complexity through high-fidelity simulations Hay et al. (2011); Herz et al. (2006); Almog and Korngreen (2016). However, they require a lot of computing resources to run and typically operate at a very fine level of granularity that does not facilitate the extraction of higher-level insights into the neuron’s computational principles. A promising approach to deriving such higher-level insights from simulations is through the training of surrogate phenomenological neuron models. These models are designed to replicate the output of high-fidelity biophysical simulations but use simplified interpretable components. This approach was employed, for example, to model computation in the dendritic tree via simple two-layer ANN Poirazi et al. (2003); Tzilivaki et al. (2019); Ujfalussy et al. (2018). Building on this line of research, a recent study by Beniaguev et al. (2021) developed a temporal convolutional network to capture the spike-level input/output (I/O) relationship with millisecond precision, accounting for the complexity of integrating diverse synaptic input across the entirety of the dendritic tree of a high-fidelity biophysical neuron model. It was found that a highly expressive temporal convolutional network with millions of parameters was essential to reproduce the aforementioned I/O relationship.
@@
In this work, we propose that a model equipped with appropriate inductive biases that align with the operational principles of a cortical neuron should be capable of capturing the I/O relationship using a substantially smaller model size. To achieve this, a model would likely need to integrate multiple mechanisms of neural expressivity and judiciously allocate computational resources and parameters in close analogy to biological neurons. Should such a construction be possible, the requisite design choices may yield insights into higher-level principles of neural computation. We proceed to design the Expressive Leaky Memory (ELM) neuron model (see Figure 1), a biologically inspired, computationally expressive, yet parameter-efficient phenomenological model of a cortical neuron. The ELM neuron functions as a recurrent cell and can be conveniently used as a drop-in replacement for LSTMs Hochreiter and Schmidhuber (1997).
@@
Our experiments show that a variant of the ELM neuron is expressive enough to accurately match the spike level I/O of a detailed biophysical model of a layer 5 pyramidal neuron at a millisecond temporal resolution with merely 8K parameters, in stark contrast to the millions of parameters required by temporal convolutional networks. To understand the implications of this design, we first probe the ELM neuron for its temporal information integration capabilities on a challenging biologically-inspired dataset requiring the addition of spike-encoded spoken digits. We find that the ELM neuron can outperform classic LSTMs, highlighting ELM neuron’s ability to leverage long-term memory. We subsequently evaluate the ELM neuron on three well-established long sequence modeling tasks prevalent in the machine learning literature, including the notoriously challenging Pathfinder-X task where it achieves $77.3\%$ accuracy even though many transformer-based models struggle to learn at all.
@@
    We propose the Expressive Leaky Memory (ELM) neuron model, a recurrent architecture that combines expressiveness with computational and parameter efficiency. The design of the ELM neuron is inspired by the structure and functionality of a biological cortical neuron.
@@
    Our experiments reveal that the ELM neuron effectively expresses the input/output relationship of a sophisticated biophysical model of a cortical neuron, suggesting that the inductive biases we introduced into the ELM neuron align with the operating principles of biological neurons.
@@
    We further demonstrate the ELM neuron’s ability to integrate temporal information over extended time horizons. Additionally, we show its capacity to handle complex long-range sequence modeling tasks that are common in machine learning, affirming its potential applicability in broader machine learning contexts.
@@ 2 The Expressive Leaky Memory Neuron
Refer to caption (b) The ELM Neuron Architecture $\begin{split}{\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\kappa_{m}}}&=\exp(\nicefrac{{-\Delta t}}{{{\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\tau_{m}}}}})\\ {\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{\kappa_{s}}}&=\exp(\nicefrac{{-\Delta t}}{{{\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{\tau_{s}}}}})\\ {\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{s_{t}}}&={\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{\kappa_{s}}}\odot{\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{s_{t-1}}}+{\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{w_{s}}}\odot\boldsymbol{x_{t}}\\ {\color[rgb]{0.87109375,0.5625,0.01953125}\boldsymbol{\Delta m_{t}}}&=\tanh(\text{MLP}_{{\color[rgb]{0.87109375,0.5625,0.01953125}\boldsymbol{w_{p}}}}([{\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{s_{t}}},{\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\kappa_{m}}}\odot{\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{m_{t-1}}}]))\\ {\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{m_{t}}}&={\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\kappa_{m}}}\odot{\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{m_{t-1}}}+{\color[rgb]{0.87109375,0.5625,0.01953125}\lambda}(1-{\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\kappa_{m}}})\odot{\color[rgb]{0.87109375,0.5625,0.01953125}\boldsymbol{\Delta m_{t}}}\\ {\color[rgb]{0.80078125,0.46875,0.73828125}\boldsymbol{y_{t}}}&={\color[rgb]{0.80078125,0.46875,0.73828125}\boldsymbol{w_{y}}}\cdot{\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{m_{t}}}\end{split}$ (1) (c) The ELM Neuron Equations
@@
Figure 1: The biologically motivated Expressive Leaky Memory (ELM) neuron model. The architecture can be divided into the following components; the input current synapse dynamics, the integration mechanism dynamics, the leaky memory dynamics, and the output dynamics. a) Sketch of a biological cortical pyramidal neuron segmented into the analogous architectural components using the corresponding colors. b) Schematics of the ELM neuron architecture, component-wise colored accordingly. c) The ELM neuron equations, where $\boldsymbol{x_{t}}\in\mathbb{R}^{d_{s}}$ is the input at time $t$, $\Delta t\in\mathbb{R}^{+}$ the fictitious elapsed time in milliseconds between two consecutive inputs $\boldsymbol{x_{t-1}}$ and $\boldsymbol{x_{t}}$ , ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{m}}\in\mathbb{R}^{d_{m}}$ are memory units, ${\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{s}}\in\mathbb{R}^{d_{s}}$ the synapse currents (traces), ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\tau_{m}}}\in\mathbb{R^{+}}^{d_{m}}$ and ${\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{\tau_{s}}}\in\mathbb{R^{+}}^{d_{s}}$ their respective timescales in milliseconds, ${\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{w_{s}}}\in\mathbb{R^{+}}^{d_{s}}$ are synapse weights, ${\color[rgb]{0.87109375,0.5625,0.01953125}\boldsymbol{w_{p}}}$ the weights of a Multilayer Perceptron (MLP) with $l_{mlp}$ hidden layers of size $d_{\mathrm{mlp}}$, ${\color[rgb]{0.80078125,0.46875,0.73828125}\boldsymbol{w_{y}}}\in\mathbb{R}^{d_{o}\times d_{m}}$ the output weights, ${\color[rgb]{0.87109375,0.5625,0.01953125}\lambda}\in\mathbb{R^{+}}$ a scaling factor for the delta memory ${\color[rgb]{0.87109375,0.5625,0.01953125}\Delta m_{t}}\in\mathbb{R}^{d_{m}}$, and ${\color[rgb]{0.80078125,0.46875,0.73828125}\boldsymbol{y}}\in\mathbb{R}^{d_{o}}$ the output.
In this section, we discuss the design of the Expressive Leaky Memory (ELM) neuron, biologically motivating the individual computational components and the overall recurrent cell architecture.
2.1 Architectural Components
@@
Neurons typically receive inputs at their synapses in the form of sparse binary events, known as spikes, which are generally converted into small positive or negative currents exhibiting characteristic temporal dynamics Kandel et al. (2000). A prevalent abstraction in computational neuroscience for this dynamic is the current-based synapse model Dayan and Abbott (2005); Burkitt (2006), which we also adopt in the Expressive Leaky Memory (ELM) neuron model (refer to the green components in Figure 1). In essence, ${\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{s_{t}}}$ represents a low-pass filtered version of potentially sparse binary input, rendering it beneficial for coincidence detection or synaptic information integration König et al. (1996). The synapse weights, ${\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{w_{s}}}$, function as a straightforward gating mechanism, serving to either amplify or attenuate the input.
@@
The state of a biological neuron can be characterized by diverse measurable quantities, such as the membrane voltage, various ion or protein concentrations (e.g. Ca+, synaptic vesicles, ATP, etc.), and their slow or fast decay (determined by their respective timescale constants). Each of these quantities carries information through time, acting as a sort of memory mechanism or hidden state Kandel et al. (2000); Dayan and Abbott (2005) (see blue components in figure 1). Most computational neuroscience models a priori fix which quantities to track (usually the membrane voltage or certain ion concentrations and their interactions) and the relationships between them (through differential equations). However, which of these biological quantities are computationally relevant, how they interact, how many integration zones a neuron has, and what the relevant timescales of these processes remain topics of active debate Aru et al. (2020); Herz et al. (2006); Almog and Korngreen (2016); Koch (1997); Chavlis and Poirazi (2021). Therefore, to match a biological neuron’s computations and expressivity accurately, the surrogate model architecture needs to accommodate a large range of possibilities. In the ELM neuron, we achieve this by making the number of memory units $d_{m}$ a hyperparameter, unlike most other computational neuroscience models. Additionally, motivated by the large diversity of timescales found within and across neurons Cavanagh et al. (2020), and their computational importance Gjorgjieva et al. (2016); Gao et al. (2020); Zeraati et al. (2023), the memory units ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{m_{t}}}$ are equipped with individually learnable memory timescales ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\tau_{m}}}$, defining their respective rate of decay (large timescales meaning slower decay), constituting a learnable yet constant forgetting mechanism.
@@
This dynamic refers to how dependent on the previous memory units ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{m_{t-1}}}$ the synaptic input ${\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{s_{t}}}$ is integrated into the memory unit updates ${\color[rgb]{0.87109375,0.5625,0.01953125}\boldsymbol{\Delta m_{t}}}$ in analogy to the dendritic tree of cortical neuron (orange components in Figure 1). We choose to parameterize this transformation using a Multilayer Perceptron (MLP), which is motivated by the ongoing discussion regarding the complexity and functional form of synaptic information integration over the dendritic tree in cortical neurons. While earlier perspectives suggested an integration process akin to linear summation Jolivet et al. (2008), newer studies advocate for complex non-linear integration Almog and Korngreen (2016); Gidon et al. (2020); Larkum (2022), specifically proposing multi-layered ANNs as a more suitable model Poirazi et al. (2003); Jadi et al. (2014); Tzilivaki et al. (2019); Marino (2021); Jones and Kording (2021, 2022); Hodassman et al. (2022), making an MLP a natural choice. Furthermore, its ability to incorporate previous memory units ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{m_{t-1}}}$ into the process accommodates state-dependent synaptic integration and related computations Hodgkin and Huxley (1952); Gasparini and Magee (2006); Bicknell and Häusser (2021), and enables the relationships between memory units ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{m}}$ to be fully learnable, irrespective of their quantity. These features provide our neuron model with enhanced flexibility over those using pre-determined differential equations.

### v3

# The Expressive Leaky Memory Neuron: an Efficient and Expressive Phenomenological Neuron Model Can Solve Long-Horizon Tasks

@@ Abstract
Biological cortical neurons are remarkably sophisticated computational devices, temporally integrating their vast synaptic input over an intricate dendritic tree, subject to complex, nonlinearly interacting internal biological processes. A recent study proposed to characterize this complexity by fitting accurate surrogate models to replicate the input-output relationship of a detailed biophysical cortical pyramidal neuron model and discovered it needed temporal convolutional networks (TCN) with millions of parameters. Requiring these many parameters, however, could stem from a misalignment between the inductive biases of the TCN and cortical neuron’s computations. In light of this, and to explore the computational implications of leaky memory units and nonlinear dendritic processing, we introduce the Expressive Leaky Memory (ELM) neuron model, a biologically inspired phenomenological model of a cortical neuron. Remarkably, by exploiting such slowly decaying memory-like hidden states and two-layered nonlinear integration of synaptic input, our ELM neuron can accurately match the aforementioned input-output relationship with under ten thousand trainable parameters. To further assess the computational ramifications of our neuron design, we evaluate it on various tasks with demanding temporal structures, including the Long Range Arena (LRA) datasets, as well as a novel neuromorphic dataset based on the Spiking Heidelberg Digits dataset (SHD-Adding). Leveraging a larger number of memory units with sufficiently long timescales, and correspondingly sophisticated synaptic integration, the ELM neuron displays substantial long-range processing capabilities, reliably outperforming the classic Transformer or Chrono-LSTM architectures on LRA, and even solving the Pathfinder-X task with over $70\%$ accuracy (16k context length). These findings raise further questions about the computational sophistication of individual cortical neurons and their role in extracting complex long-range temporal dependencies.
@@ 1 Introduction
The human brain has impressive computational capabilities, yet the precise mechanisms underpinning them remain largely undetermined. Two complementary directions are pursued in search of mechanisms for brain computations. On the one hand, many researchers investigate how these capabilities could arise from the collective activity of neurons connected into a complex network structure Maass (1997); Gerstner & Kistler (2002); Grüning & Bohte (2014), where individual neurons might be as basic as leaky integrators or ReLU neurons. On the other hand, it has been proposed that the intrinsic computational power possessed by individual neurons Koch (1997); Koch & Segev (2000); Silver (2010) contributes a significant part to the computations.
@@
Even though most work focuses on the former hypothesis, an increasing amount of evidence indicates that cortical neurons are remarkably sophisticated Silver (2010); Gidon et al. (2020); Larkum (2022), even comparable to expressive multilayered artificial neural networks Poirazi et al. (2003); Jadi et al. (2014); Beniaguev et al. (2021); Jones & Kording (2021), and capable of discriminating between dozens to hundreds of input patterns Gütig & Sompolinsky (2006); Hawkins & Ahmad (2016); Moldwin & Segev (2020). Numerous biological mechanisms, such as complex ion channel dynamics (e.g. NMDA nonlinearity Major et al. (2013); Lafourcade et al. (2022); Tang et al. (2023)), plasticity on various and especially longer timescales (e.g. slow spike frequency adaptation Kobayashi et al. (2009); Bellec et al. (2018)), the intricate cell morphology (e.g. nonlinear integration by dendritic tree Stuart & Spruston (2015); Poirazi & Papoutsi (2020); Larkum (2022)), and their interactions, have been identified to contribute to their complexity.
@@
Detailed biophysical models of cortical neurons aim to capture this inherent complexity through high-fidelity mechanistic simulations Hay et al. (2011); Herz et al. (2006); Almog & Korngreen (2016). However, they require a lot of computing resources to run and typically operate at a very fine level of granularity that does not facilitate the extraction of higher-level insights into the neuron’s computational principles. A promising approach to derive such higher-level insights from simulations is through the training of surrogate phenomenological neuron models. Such models are designed to replicate the output of biophysical simulations but use simplified interpretable components. This approach was employed, for example, to model computation in the dendritic tree via simple two-layer ANN Poirazi et al. (2003); Tzilivaki et al. (2019); Ujfalussy et al. (2018). Building on this line of research, a recent study by Beniaguev et al. (2021) developed a temporal convolutional network to capture the spike-level input/output (I/O) relationship with millisecond precision, accounting for the complexity of integrating diverse synaptic input across the entirety of the dendritic tree of a high-fidelity biophysical neuron model. It was found that a highly expressive temporal convolutional network with millions of parameters was essential to reproduce the aforementioned I/O relationship.
@@
In this work, we propose that a model equipped with appropriate biologically inspired components that align with the high-level computational principles of a cortical neuron should be capable of capturing the I/O relationship using a substantially smaller model size. To achieve this, a model would likely need to account for multiple mechanisms of neural expressivity and judiciously allocate computational resources and parameters in a rough analogy to biological neurons. Should such a construction be possible, the required design choices may yield insights into principles of neural computation at the conceptual level. We proceed to design the Expressive Leaky Memory (ELM) neuron model (see Figure 1), a biologically inspired phenomenological model of a cortical neuron. While biologically inspired, low-level biological processes are abstracted away for computational efficiency, and consequently, individual parameters of the ELM neuron are not designed for direct biophysical interpretability. Nevertheless, model ablations can provide conceptual insights into the computational components required to emulate the cortical input/output relationship. The ELM neuron functions as a recurrent cell and can be conveniently used as a drop-in replacement for LSTMs Hochreiter & Schmidhuber (1997).
@@
Our experiments show that a variant of the ELM neuron is expressive enough to accurately match the spike level I/O of a detailed biophysical model of a layer 5 pyramidal neuron at a millisecond temporal resolution with a few thousand parameters, in stark contrast to the millions of parameters required by temporal convolutional networks. Conceptually, we find accurate surrogate models to require multiple memory-like hidden states with longer timescales and highly nonlinear synaptic integration. To explore the implications of neuron-internal timescales and sophisticated synaptic integration into multiple memory units, we first probe its temporal information integration capabilities on a challenging biologically inspired neuromorphic dataset requiring the addition of spike-encoded spoken digits. We find that the ELM neuron can outperform classic LSTMs leveraging a sufficient number of slowly decaying memory and highly nonlinear synaptic integration. We subsequently evaluate the ELM neuron on the well-established long sequence modeling LRA benchmarks from the machine learning literature, including the notoriously challenging Pathfinder-X task, where it achieves over $70\%$ accuracy but many transformer-based models do not learn at all.
@@
    We propose the phenomenological Expressive Leaky Memory (ELM) neuron model, a recurrent cell architecture inspired by biological cortical neurons.
@@
    The ELM neuron efficiently learns the input/output relationship of a sophisticated biophysical model of a cortical neuron, indicating its inductive biases to be well aligned.
@@
    The ELM neuron facilitates the formulation and validation of hypotheses regarding the underlying high-level neuronal computations using suitable architectural ablations.
  4. 4.
    Lastly, we demonstrate the considerable long-sequence processing capabilities of the ELM neuron through the use of long memory and synapse timescales.
@@ 2 The Expressive Leaky Memory Neuron
Refer to caption (b) The ELM Neuron Architecture $\begin{split}{\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\kappa_{m}}}&=\exp(\nicefrac{{-\Delta t}}{{{\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\tau_{m}}}}})\\ {\color[rgb]{0.0078125,0.62109375,0.44921875}\bm{\kappa_{s}}}&=\exp(\nicefrac{{-\Delta t}}{{{\color[rgb]{0.0078125,0.62109375,0.44921875}\bm{\tau_{s}}}}})\\ {\color[rgb]{0.0078125,0.62109375,0.44921875}\bm{s_{t}}}&={\color[rgb]{0.0078125,0.62109375,0.44921875}\bm{\kappa_{s}}}\odot{\color[rgb]{0.0078125,0.62109375,0.44921875}\bm{s_{t-1}}}+{\color[rgb]{0.0078125,0.62109375,0.44921875}\bm{w_{s}}}\odot\bm{x_{t}}\\ {\color[rgb]{0.87109375,0.5625,0.01953125}\bm{\Delta m_{t}}}&=\tanh(\text{MLP}_{{\color[rgb]{0.87109375,0.5625,0.01953125}\bm{w_{p}}}}([{\color[rgb]{0.0078125,0.62109375,0.44921875}\bm{s_{t}}},{\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\kappa_{m}}}\odot{\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{m_{t-1}}}]))\\ {\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{m_{t}}}&={\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\kappa_{m}}}\odot{\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{m_{t-1}}}+{\color[rgb]{0.87109375,0.5625,0.01953125}\lambda}\cdot(1-{\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\kappa_{m}}})\odot{\color[rgb]{0.87109375,0.5625,0.01953125}\bm{\Delta m_{t}}}\\ {\color[rgb]{0.80078125,0.46875,0.73828125}\bm{y_{t}}}&={\color[rgb]{0.80078125,0.46875,0.73828125}\bm{w_{y}}}\cdot{\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{m_{t}}}\end{split}$ (1) (c) The ELM Neuron Equations
@@
Figure 1: The biologically motivated Expressive Leaky Memory (ELM) neuron model. The architecture can be divided into the following components: the input current synapse dynamics, the integration mechanism dynamics, the leaky memory dynamics, and the output dynamics. a) Sketch of a biological cortical pyramidal neuron segmented into the analogous architectural components using the corresponding colors. b) Schematics of the ELM neuron architecture, component-wise colored accordingly. c) The ELM neuron equations, where $\bm{x_{t}}\in\mathbb{R}^{d_{s}}$ is the input at time $t$, $\Delta t\in\mathbb{R}^{+}$ the fictitious elapsed time in milliseconds between two consecutive inputs $\bm{x_{t-1}}$ and $\bm{x_{t}}$ , ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{m}}\in\mathbb{R}^{d_{m}}$ are memory units, ${\color[rgb]{0.0078125,0.62109375,0.44921875}\bm{s}}\in\mathbb{R}^{d_{s}}$ the synapse currents (traces), ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\tau_{m}}}\in\mathbb{R^{+}}^{d_{m}}$ and ${\color[rgb]{0.0078125,0.62109375,0.44921875}\bm{\tau_{s}}}\in\mathbb{R^{+}}^{d_{s}}$ their respective timescales in milliseconds, ${\color[rgb]{0.0078125,0.62109375,0.44921875}\bm{w_{s}}}\in\mathbb{R^{+}}^{d_{s}}$ are synapse weights, ${\color[rgb]{0.87109375,0.5625,0.01953125}\bm{w_{p}}}$ the weights of a Multilayer Perceptron (MLP) with $l_{\mathrm{mlp}}$ hidden layers of size $d_{\mathrm{mlp}}$, ${\color[rgb]{0.80078125,0.46875,0.73828125}\bm{w_{y}}}\in\mathbb{R}^{d_{o}\times d_{m}}$ the output weights, ${\color[rgb]{0.87109375,0.5625,0.01953125}\lambda}\in\mathbb{R^{+}}$ a scaling factor for the delta memory ${\color[rgb]{0.87109375,0.5625,0.01953125}\Delta m_{t}}\in\mathbb{R}^{d_{m}}$, and ${\color[rgb]{0.80078125,0.46875,0.73828125}\bm{y}}\in\mathbb{R}^{d_{o}}$ the output.
In this section, we discuss the design of the Expressive Leaky Memory (ELM) neuron, and its variant Branch-ELM. Its architecture is engineered to capture sophisticated cortical neuron computations efficiently. Abstracting mechanistic neuronal implementation details away, we resort to an overall recurrent cell architecture with biologically motivated computational components. This design approach emphasizes conceptual over mechanistic insight into cortical neuron computations.
@@
Neurons receive inputs at their synapses in the form of sparse binary events known as spikes Kandel et al. (2000). While the Excitatory/Inhibitory synapse identity determines the sign of the input (always given), the positive synapse weights can act as simple input gating ${\color[rgb]{0.0078125,0.62109375,0.44921875}\bm{w_{s}}}$ (learned in Branch-ELM). The synaptic trace ${\color[rgb]{0.0078125,0.62109375,0.44921875}\bm{s_{t}}}$ denotes a filtered version of the input, believed to aid coincidence detection and synaptic information integration in neurons König et al. (1996). This implementation is known as the current-based synapse dynamic Dayan & Abbott (2005).
@@
The state of a biological neuron may be characterized by diverse measurable quantities, such as their membrane voltage or various ion/molecule concentrations (e.g. Ca+, mRNA, etc.), and their rate of decay over time (slow decay <-> large timescale), endowing them with a sort of leaky memory Kandel et al. (2000); Dayan & Abbott (2005). However, which of these quantities are computationally relevant, how and where they interact, and on what timescale, remains a topic of active debate Aru et al. (2020); Herz et al. (2006); Almog & Korngreen (2016); Koch (1997); Chavlis & Poirazi (2021); Cavanagh et al. (2020); Gjorgjieva et al. (2016). Therefore, to match a biological neuron’s computations, the surrogate model architecture needs to be expressive enough to accommodate a large range of possibilities. In the ELM neuron, we achieve this by making the number of memory units $d_{m}$ a hyper-parameter and equipping each of them with a ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{\tau_{m}}}$ (always learnable), setting it apart most other computational neuroscience models.
@@
This dynamic refers to how the synaptic input ${\color[rgb]{0.0078125,0.62109375,0.44921875}\bm{s_{t}}}$ is integrated into the memory units ${\color[rgb]{0.87109375,0.5625,0.01953125}\bm{\Delta m_{t}}}$ in analogy to the dendritic tree of a cortical neuron. While earlier perspectives suggested an integration process akin to linear summation Jolivet et al. (2008), newer studies advocate for complex nonlinear integration Almog & Korngreen (2016); Gidon et al. (2020); Larkum (2022), specifically proposing multi-layered ANNs as more suitable descriptions Poirazi et al. (2003); Jadi et al. (2014); Marino (2021); Jones & Kording (2021); Iatropoulos et al. (2022); Jones & Kording (2022); Hodassman et al. (2022), also backed by recent evidence of neuronal plasticity beyond synapses Losonczy et al. (2008); Holtmaat et al. (2009); Abraham et al. (2019). Motivated by this ongoing discussion, we choose to parameterize the input integration using a Multilayer Perceptron (MLP) (${\color[rgb]{0.87109375,0.5625,0.01953125}\bm{w_{p}}}$ always learnable, with $l_{\mathrm{mlp}}=1$ and $d_{\mathrm{mlp}}=2d_{m}$), which can be used to explore the full range of hypothesized integration complexities, while offering a straightforward way to quantify and ablate the ELM neuron integration complexity. In the Branch-ELM variant (for motivation see details in Section 4 and Figure 4) we extend the integration mechanism dynamics; before the MLP is applied, the synaptic input ${\color[rgb]{0.0078125,0.62109375,0.44921875}\bm{s_{t}}}\in\mathbb{R}^{d_{s}}$ is reduced to a smaller number ($d_{\mathrm{tree}}$) of branch activations, each computed as a sum over $d_{\mathrm{brch}}$ neighboring synaptic inputs (with $d_{\mathrm{tree}}*d_{\mathrm{brch}}=d_{s}$). In this variant the ${\color[rgb]{0.0078125,0.62109375,0.44921875}\bm{w_{s}}}$ need to be learnable, as they are responsible for weighting the sum and cannot be absorbed in the MLP later. Despite the biological inspiration, the MLP and synapses are only intended to capture the neuron analogous plasticity and dendritic nonlinearity, and cannot give a mechanistic explanation of these phenomena in neuron. Finally, incorporating previous memory units ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{m_{t-1}}}$ into the integration process, the ELM can accommodate state-dependent synaptic integration and related computations Hodgkin & Huxley (1952); Gasparini & Magee (2006); Bicknell & Häusser (2021), and enables the relationships among memory units ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{m}}$ to be fully learnable. The range of the ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{m}}$ values is controlled by ${\color[rgb]{0.87109375,0.5625,0.01953125}\lambda}$, and the mixing of the proposal values by the parameter ${\color[rgb]{0.00390625,0.44921875,0.69921875}\bm{k_{m}}}$ (for details on parameters see Appendix Section A). Crucially, our approach sidesteps the need for expert-designed and pre-determined differential equations typical in phenomenological neuron modeling.